3. Complete the following table (fill all the empty cells): (1)

Total output Fixed cost Variable cost Total cost Marginal cost Average fixed cost Average variable cost Average total cost
0 100 $0 100 $------ $------ $------ $------
1 100 $10
2 100 $18 118 $5
3 100 $23 123
4 100 $6.50
1
Expert's answer
2020-02-26T10:00:38-0500

Q FC VC TC MC AFC AVC ATC

0 100 $0 100 - - - -

1 100 $10 110 10 100 10 11

2 100 $18 118 8 50 9 59

3 100 $23 123 5 33.3 7.67 41

4 100 $26 126 3 25 6.5 31.5

TC = VC + FC,

MC = change in TC / change in Q,

AFC = FC/Q, AVC = VC/Q, ATC = TC/Q.